OpenStudy (anonymous):

Larry has taken out a loan for college. He started paying off the loan with a first payment of $150. Each month he pays, he wants to pay back 1.3 times as the amount he paid the month before. Explain to Larry how to represent his first 15 payments in sigma notation. Then explain how to find the sum of his first 15 payments, using complete sentences. Explain why this series is convergent or divergent.

OpenStudy (anonymous):

@Michele_Laino can you help again please

OpenStudy (michele_laino):

please wait a moment

OpenStudy (michele_laino):

the first payment is 150 the second payment is 150-1.3*150 the third payment is 150-1.3*(150-1.3*150) and so on...

OpenStudy (anonymous):

I think there should be a faster way because there is 15 payments

OpenStudy (michele_laino):

yes! I wrote those formulas in order to show the pattern

OpenStudy (michele_laino):

I call with a the quantity 150, namely a= 150, then I call with k the percentage, namely k=1.3

OpenStudy (michele_laino):

so I can write this first payment= a second payment =a-ka third payment= a-k(a-ka)=a-ka-k^2 a fourth payment= a-k(a-ka-k^2 a)= a-ka -k^2 a- k^3 a

OpenStudy (michele_laino):

n-th payment= a-ka-k^a-...-k^(n-1) a

OpenStudy (michele_laino):

oops..n-th payment= a-ka-k^2 a-...-k^(n-1) a

OpenStudy (anonymous):

I'm confused and a bit discouraged

OpenStudy (michele_laino):

I have written, using formulas, the text of your nproblem

OpenStudy (anonymous):

How would I explain this using steps

OpenStudy (michele_laino):

here is the sigma-notation for n-th payment p(n): \[\Large p\left( n \right) = a - a\sum\limits_{i = 1}^{n - 1} {{k^i}} \]

OpenStudy (michele_laino):

you can explain, using my replies above

OpenStudy (anonymous):

oh ok

OpenStudy (michele_laino):

my for mula above, namely: \[\Large p\left( n \right) = a - a\sum\limits_{i = 1}^{n - 1} {{k^i}} \] is valid for n>1 whereas for n=1, we have p(1)= a=150

OpenStudy (michele_laino):

formula*

OpenStudy (anonymous):

ok hold on

OpenStudy (michele_laino):

ok! please wait...

OpenStudy (anonymous):

ok i got it now

OpenStudy (michele_laino):

now we have to compute this sum: S=p(1)+p(2)+...+p(15) so we can write this: \[\Large \begin{gathered} S = p\left( 1 \right) + p\left( 2 \right) + p\left( 3 \right) + ... + p\left( {15} \right) = \hfill \\ = a + \sum\limits_{j = 2}^{15} {p\left( j \right)} = a + \sum\limits_{j = 2}^{15} {\left( {a - a\sum\limits_{i = 1}^{j - 1} {{k^i}} } \right)} \hfill \\ \end{gathered} \]